# Noemi Walczak Université Paris 1 Panthéon-Sorbonne, Paris, France Kenan Huremović Armando Rungi

### v3

@@ 5. RESULTS
where $\bm{\tilde{\rho}}=diag(\tilde{\rho}_{1}^{1},...,\tilde{\rho}_{N}^{J})\in\mathcal{M}(NJ\times NJ)$ is a diagonal matrix with entries $\tilde{\rho}_{i}^{j}\equiv e_{i}^{j}/p_{i}^{j}q_{i}^{j}\ \forall i\in\mathcal{N},j\in\mathcal{J}$ corresponding to the tons of emissions per dollar of output in a given sector-country pair, $\mathbf{\tilde{\Omega}}_{EU}\in\mathcal{M}(NJ\times NJ)$ is the matrix of input-output coefficients, with non-zero entries only for European destination sectors, excluding intra-EU trade, $\mathbf{\Lambda}\in\mathbb{R}^{NJ}$ is the vector of Domar weights and $GNE$ is the world Gross National Expenditure. The term $\gamma\mathbf{\tilde{\Omega}}_{EU}diag(\mathbf{\Lambda})\mathbf{1}GNE$ corresponds to the total value of European imports, expressed compactly as $\mathbf{X}_{EU}\mathbf{1}$, where $\mathbf{X}_{EU}\in\mathcal{M}(NJ\times NJ)$ is the matrix of trade flows between all country-sector pairs, with non-zero entries only for European destination sectors, again excluding intra-EU trade. The expression captures the total emissions embodied in European imports, including both emissions generated from the production of goods exported directly to the EU (emissions embodied in direct imports) and emissions generated throughout the supply chain (emissions from indirect imports), accounted for by the Leontief inverse $(\mathbf{I}-\mathbf{\Omega})^{-1}$. If we abstract from the upstream linkages by setting $\mathbf{\Omega}=\mathbf{0}$, only emissions embodied in direct imports remain, reflecting emissions from goods produced abroad and directly exported to the EU.
Table 2 presents the changes in emissions associated with European imports. Our results reported in Panel (a) show that CBAM leads to a 4.73% reduction in emissions embodied in direct imports, reaching almost 7% when only dirty intermediate goods are considered. When the indirect emissions are incorporated (Panel (b)), the magnitude of emissions reductions is attenuated but remains substantial: total emissions reduction under the current CBAM is 2.96%, compared to 5.19% under the full CBAM. This attenuation is expected. While the CBAM directly reduces imports of carbon-intensive goods explicitly covered by the adjustment mechanism – captured by the reduction in emissions embodied in direct imports – it simultaneously induces substitution towards non-targeted inputs. This effect propagates upstream through the supply network, increasing the demand for inputs along the supply chain, some of which are dirty, thereby raising indirect emissions. These upstream emissions fall outside the scope of the CBAM. As a consequence, the observed reduction in the carbon content of imports is smaller when these indirect channels are taken into account, which is a novel insight provided by our paper. Therefore, failing to account for these higher-order network effects leads to a significant overestimation of the effects of the CBAM.
Finally, Panel (c) reports changes in emissions leakage. In our framework, the leakage coincides with the change in emissions of countries that have not adopted a carbon market yet. Emissions leakage declines by 0.07% under the current CBAM and by 0.19% under the full CBAM. This confirms that carbon border adjustments contribute to containing emissions in countries with less stringent environmental regulations. Overall, our findings suggest that the policy contributes to the reduction of emissions leakage.
Variable Current CBAM Full CBAM
@@
Table 2: Policy-induced changes in Emissions Embodied in EU Imports and Emissions Leakage (% of baseline level). Each value is considered under two scenarios: one in which the CBAM is determined endogenously (columns labeled Endog.) and one in which it is exogenous (columns labeled Exog.). Our main analysis focuses on the endogenous scenario (columns Endog.).
@@
We are also able to decompose the total change in emissions embodied in European imports into two key components: a technology effect, i.e., capturing changes in emissions from production, and a reallocation effect, i.e., capturing changes in sourcing patterns across countries and sectors. A detailed derivation of this decomposition, expressed in terms of changes in cost shares and production-based emissions is provided in A.2. For our current purposes, it is sufficient to note that the technology effect corresponds to the changes in the vector $\bm{\tilde{\rho}}$ of emissions per dollar of output in equation (17), holding the structure of global production linkages constant. Unlike $\rho_{i}^{j}=t_{i}e_{i}^{j}/p_{i}^{j}q_{i}^{j}$ that is a fixed parameter of our model, $\tilde{\rho}_{i}^{j}=e_{i}^{j}/p_{i}^{j}q_{i}^{j}$ varies in response to changes in the input mix. Intuitively, this captures the reduction in the tons of emissions generated from production resulting from a shift toward cleaner inputs within existing production chains. By contrast, the reallocation effect stems from changes in the production network, captured by the Leontief inverse $(\mathbf{I}-\mathbf{\Omega})^{-1}$, and in the European imports, represented by $\mathbf{X}_{EU}$, holding production-based emissions fixed.
@@
The corresponding Figure 4 illustrates this decomposition clearly. In all cases – whether total, clean, or dirty imports, and under both the current and full CBAM scenarios – the reallocation effect accounts for more than 50% of the total change in emissions. This evidence underscores the dominant role of trade reallocation mechanisms and highlights the importance of tracking changes that occur across the production network in order to understand the environmental impact of border carbon adjustments. In conclusion, our results underscore the importance of adopting a supply-chain-wide perspective in both the evaluation and design of border carbon adjustments, thus emphasising the added value of an endogenous CBAM mechanism to ensure that its full emissions footprint is captured.
@@
We close this section by examining the effects of the CBAM on real GNE. Table 3 reports the percentage changes in EU (Panel (a)) and extra-EU (Panel (b)) real gross national expenditures following the introduction of the CBAM. Under the current CBAM configuration, EU GNE slightly rises by 0.006%, while under the full CBAM scenario, the increase reaches 0.04%. This modest improvement in the EU GNE is driven by the reallocation of demand away from foreign carbon-intensive goods and towards domestic production. Given that trade deficits are fixed in the model, this translates into an increase in the price of domestic goods relative to foreign ones. This effect improves the EU’s trade balance and raises GNE. Results are consistent with [19], who report a welfare gain of 0.016%. Despite the increase in GNE, real wages decline by 0.02% and 0.05% under the current and full CBAM scenarios, respectively. Panel (b) considers the effects on GNE in extra-EU countries. In these economies, the impact on GNE is modestly negative, with decreases of 0.008% under the current CBAM and 0.02% under the full CBAM scenario. Notably, the endogenous adjustment of the CBAM does not substantially alter the magnitude of the changes in real GNE and wages, which remain close to the scenario where the CBAM is exogenous.
Variable Current CBAM Full CBAM
@@
Table 3: Policy-induced changes in Gross National Expenditure (% of baseline level). Each value is considered under two scenarios: one in which the CBAM is determined endogenously (columns labeled Endog.) and one in which it is exogenous (columns labeled Exog.). Our main analysis focuses on the endogenous scenario (columns Endog.).
The role of integration and technology.
@@
Intuitively, the effect of the CBAM strongly depends both on the technology used in importing source countries as well as the integration of the EU in the global supply network. We now conduct two kinds of counterfactual exercises to shed some light on how these two dimensions influence the effects of the CBAM. In the first, to capture the effect of technology, we vary the output elasticity with respect to emissions, while in the second, we vary the integration of the EU in the global supply chain. In both types of exercises, we focus on the key outcome of our analysis — embodied emissions in EU imports.
First, to examine the effect of technology, we construct a counterfactual in which carbon intensities, captured with $\rho_{i}^{j}$, are halved relative to those observed at the baseline. The resulting scenario enables us to assess how the effects of CBAM respond to a counterfactual circumstance where the world’s technology becomes cleaner.
@@
Figure 5 shows the reduction in emissions embodied in EU imports—both direct and total—under two scenarios: the one with observed carbon intensities and a counterfactual in which the carbon intensity of production is halved across all country-sector pairs. Results are presented as percentage changes relative to the baseline. As expected, the impact of the CBAM on reducing embodied emissions weakens in a world where global production is less carbon-intensive, since the associated CBAM charges are lower. However, the reduction in effectiveness is more than proportional, highlighting the disproportionate influence of production technologies on the policy’s impact.
@@
Figure 5: Policy-induced changes in Emissions Embodied in EU Imports under Current and Low-Carbon Intensity Scenarios (Percentages are computed with reference to the value of the corresponding embodied emissions (direct or total) under the scenario with current carbon intensities and no CBAM applied).
Second, to assess the role of global supply chain integration, we conduct a counterfactual exercise that systematically alters the reliance on imported inputs. Specifically, for each EU importing sector, we vary the ratio of imported to domestic inputs, where a higher ratio reflects deeper integration of EU production into global supply chains. We then plot the impact of the CBAM on embodied emissions as a function of this ratio. This exercise is, therefore, informative about how the CBAM works under different levels of integration, while holding constant the relative sourcing pattern across foreign trade partners and sectors.
Figure 6 shows how emissions embodied in imports change following the introduction of the CBAM, relative to the baseline, as the degree of integration into global supply chains varies. The CBAM induces two types of substitution: towards domestic inputs and towards cleaner and untargeted foreign inputs. We find that when the EU economy is less integrated – that is, when sectors rely more on domestic rather than imported inputs – the CBAM leads to a stronger shift toward domestic sourcing. This reallocation results in a larger reduction in emissions embodied in imports, simply because a greater share of inputs originates within the EU and thus lies outside the CBAM’s scope. As integration rises, the dependence on foreign inputs increases, diminishing the potential for substituting toward domestic inputs, and consequently weakening the emissions-reducing effect of the policy. Nevertheless, the CBAM continues to influence sourcing decisions, prompting a shift toward cleaner or exempt foreign goods, which partly offsets the reduced scope for domestic substitution. Although emissions embodied in imports still fall relative to the baseline, the overall reduction is smaller, highlighting the dominant role of domestic substitution in driving the CBAM’s impact. Notably, the effect of integration on CBAM effectiveness is less than proportional.
@@
Figure 6: Policy-induced changes in Emissions Embodied in EU Imports under varying levels of trade integration (Percentages are computed with reference to the value of the corresponding embodied emissions (direct or total) under the scenario with the current level of trade integration (0% change) and no CBAM applied).
Our counterfactual exercises therefore show that cleaner global production reduces the CBAM’s impact in a more-than-proportional way, underscoring the importance of the technology dimension. A higher supply chain integration weakens the policy’s effectiveness, as the scope for substituting toward domestic inputs shrinks. Still, the integration in the global supply chain affects the effects of CBAM, albeit to a lesser extent than the technology dimension. We, however, note that these two effects are interrelated, as the EU producers in general tend to source more from destinations using more carbon-intensive technologies.
@@ 6. CONCLUSION
This paper offers a comprehensive assessment of the economic and environmental implications of the EU’s CBAM with a multi-country, multi-sector general equilibrium model with global input-output linkages. By endogenizing both the CBAM and ETS allowances’ prices, our model captures the feedback mechanisms that arise between national climate policies and international trade patterns. Results show that CBAM induces a reallocation of purchases towards domestic and foreign clean inputs and significantly reduces emissions embodied in imports. Crucially, our findings highlight the importance of accounting for supply chain reallocation effects: while the CBAM substantially reduces direct emissions embodied in imports, a non-negligible share of emissions comes from upstream inputs and remains outside the policy scope. Overall, the paper contributes to the growing literature on the problem of international coordination by open economies in the presence of climate change. In particular, our framework offers a valuable tool for policymakers to design across-the-border policies that account for the complex interdependencies of global supply chains.
@@ A.1 Data
  • Input-Output Tables To calibrate the model on the year 2018, we use the 2021 release of the OECD Inter-Country Input-Output tables [10], collecting data on expenditures in intermediate inputs $X_{ni}^{kj}$ and final goods $F_{ni}^{j}$, gross output $GO_{i}^{j}$ and value-added $VA_{i}^{j}$. Concerning final consumption, we ignore Gross Fixed Capital Formation and Changes in Inventories and Valuables. Values are reported in millions of U.S. dollars at current prices and the flows of goods and services within and across countries are directly mapped to the cost shares in intermediate goods $\tilde{\omega}_{ni}^{kj}=X_{ni}^{kj}/\sum_{n\in\mathcal{N}}\sum_{j\in\mathcal{J}}X_{ni}^{kj}$ and the shares of final consumption $\alpha_{i}^{j}=\sum_{n\in\mathcal{N}}F_{ni}^{j}/\sum_{n\in\mathcal{N}}\sum_{j\in\mathcal{J}}F_{ni}^{j}$. Finally, the shares of value added are computed as $\beta_{i}^{j}=VA_{i}^{j}/GO_{i}^{j}$. We have a unique country-sector pair (Chile - Motor vehicles, trailers and semi-trailers) with a gross output equal to 0, which can be either due to missing observations or zero production in that sector and country. Since these cases generate computational issues, we input a value of 1 to this observation.
@@
  • Emissions and carbon prices Emissions from production are taken from the OECD Environmental Statistics database. We consider Scope 1 Emissions for each country-sector pair $E_{i}^{j}$, including all direct GHG emissions measured in million tonnes of $\text{CO}_{2}$ equivalent. For the European Union, Iceland and Norway we employ a distinct approach. Emissions data are taken from the European Union Transaction Log (EUTL), recording all transactions carried out under the EU-ETS, with information on each entity covered by the system. The EUTL provides granula, entity-level information on verified emissions and the volume of freely allocated allowances on an annual basis. We aggregate these data at the selected sectoral level and adjust the reported emissions by deducting those covered by freely allocated allowances. In order to build the emissions elasticity $\rho_{i}^{j}$, these values are multiplied by the corresponding country’s Effective Carbon Rate (ECR), converted into USD using the OECD exchange rate for the year 2018. Effective Carbon Rates represent the total price $t_{i}$ of one ton of $\text{CO}_{2}$ emissions as the sum of carbon taxes, specific taxes on energy use and the price of tradable emission permits. Emission elasticities are then computed as $\rho_{i}^{j}=t_{i}E_{i}^{j}/GO_{i}^{j}$.

### v4

@@ 6. Results
Vector $\mathbf{EEI}(EU)\in\mathbb{R}^{NJ}$ has entries corresponding to the emissions embodied in European imports, as they are generated from the production of each good $j\in\mathcal{J}$ in every country $i\in\mathcal{N}$ embodied in European imports. In particular, the first equivalence in eq. (17) corresponds to the OECD definition, while the subsequent equality restates it in terms of our theoretical framework. The term $\bm{\tilde{\rho}}=diag(\tilde{\rho}_{1}^{1},...,\tilde{\rho}_{N}^{J})\in\mathcal{M}(NJ\times NJ)$ is a diagonal matrix with entries $\tilde{\rho}_{i}^{j}\equiv e_{i}^{j}/p_{i}^{j}q_{i}^{j}\ \forall i\in\mathcal{N},j\in\mathcal{J}$ corresponding to the tons of emissions per dollar of output in a given sector-country pair. The matrix $\mathbf{\tilde{\Omega}}_{EU}\in\mathcal{M}(NJ\times NJ)$ contains input-output coefficients, with non-zero entries only for European destination sectors, excluding intra-EU trade. The vector $\mathbf{\Lambda}\in\mathbb{R}^{NJ}$ contains Domar weights, and $GNE$ is the world Gross National Expenditure. The term $\gamma\mathbf{\tilde{\Omega}}_{EU}diag(\mathbf{\Lambda})\mathbf{1}GNE$ corresponds to the total value of European imports, expressed compactly as $\mathbf{X}_{EU}\mathbf{1}$, where $\mathbf{X}_{EU}\in\mathcal{M}(NJ\times NJ)$ is the matrix of trade flows between all country-sector pairs, with non-zero entries only for European destination sectors, once again excluding intra-EU trade. A more detailed derivation is left to the corresponding section in the Appendix A.2.
Briefly, The expression captures the total emissions embodied in European imports, including both emissions generated from the production of goods exported directly to the EU (emissions embodied in direct imports) and emissions generated throughout the supply chain (emissions from indirect imports), accounted for by the Leontief inverse $(\mathbf{I}-\mathbf{\Omega})^{-1}$. If we abstract from the upstream linkages by setting $\mathbf{\Omega}=\mathbf{0}$, only emissions embodied in direct imports remain, reflecting emissions from goods produced abroad and directly exported to the EU.
Variable Reduced CBAM Full CBAM
@@
Table 2: Changes in emissions embodied in EU imports, and emissions leakage (% of baseline year 2024). Each value is considered under two scenarios: one in which the CBAM is determined endogenously (Endog.) and one in which it is exogenous (Exog.)
Table 2 presents the changes in emissions associated with European imports. Our results in panel (a) show that full CBAM leads to an 8.84% reduction in emissions embodied in direct imports, reaching 13.12% when only dirty intermediate goods are considered. When the indirect emissions are incorporated (panel (b)), the magnitude of reductions is attenuated but remains substantial: total emissions under the full CBAM reduce by 5.19%. This attenuation is expected. In fact, while the CBAM directly reduces imports of carbon-intensive goods explicitly covered by the adjustment mechanism – captured by the reduction in emissions embodied in direct imports – it simultaneously induces substitution towards non-targeted inputs. This effect propagates upstream through the supply network, increasing the demand for upstream inputs, some of which are dirty, thereby raising indirect emissions. These upstream emissions are not accounted for in the design of CBAM. Briefly, failing to account for higher-order network effects results in a significant overestimation of CBAM efficacy, which is a novel result of our paper. As shown in Appendix Table A4, a higher elasticity amplifies the effects.
Finally, panel (c) reports changes in emissions leakage. In our framework, the leakage coincides with the change in emissions of countries that have not yet adopted a national carbon market. Emissions leakage declines by 0.07% under the reduced CBAM and by 0.19% under the full CBAM. Thus, leakage still exists after CBAM, and it is only partially impacted by the border adjustment mechanism. See also the comments in Section 6.1. Given the current institutional setting, we are unable to observe the strategic responses to the CBAM, which can help converge on policy targets.
Notably, we propose a decomposition of the total change in emissions embodied in European imports into two key components:
  1. 1.
    a technology effect, which captures changes in emissions from production,
  2. 2.
    a reallocation effect, which captures changes in sourcing patterns across countries and sectors.
@@
A detailed derivation of the decomposition is provided in A.2. For our current purposes, it is sufficient to note that the technology effect corresponds to the changes in the vector $\bm{\tilde{\rho}}$ of emissions per dollar of output in equation (17), holding the structure of global production linkages constant. Unlike $\rho_{i}^{j}=t_{i}e_{i}^{j}/p_{i}^{j}q_{i}^{j}$, which is a fixed parameter of our model, $\tilde{\rho}_{i}^{j}=e_{i}^{j}/p_{i}^{j}q_{i}^{j}$ varies in response to changes in the input mix. Intuitively, this captures the reduction in the tons of emissions generated from production resulting from a shift toward cleaner inputs, given existing production chains. By contrast, the reallocation effect stems from changes in the production network, captured by the Leontief inverse $(\mathbf{I}-\mathbf{\Omega})^{-1}$, and in the European imports, represented by $\mathbf{X}_{EU}$, holding production of emissions fixed.
@@
The corresponding Figure 4 illustrates the results of the decomposition. In all cases – whether total, clean, or dirty imports, and under both the reduced and full CBAM scenarios – the reallocation effect accounts for more than 50% of the total change in emissions. This evidence underscores the dominant role of trade reallocation mechanisms and highlights the importance of tracking changes that occur across the production network in order to understand the environmental impact of carbon border adjustments. In conclusion, we argue, it is important to adopt a supply-chain-wise perspective to ensure that full emissions’ footprint is captured.
@@
6.3 Welfare gains.
In the following paragraphs, we examine the effects of the CBAM on real GNE, as the latter proxy our welfare gains [^16] Please note how welfare gains are driven by assumptions, among others, on fixed deficits and inelastic labor supply.. Table 3 reports the percentage changes in EU (panel (a)) and extra-EU (panel (b)) real gross national expenditures following the introduction of the CBAM. Under the reduced configuration, EU GNE slightly increases by 0.006%, while under the full CBAM scenario, the increase reaches 0.04%. This modest improvement in the EU GNE is driven by the reallocation of demand away from foreign carbon-intensive goods and toward domestic production. Given that trade deficits are fixed in the model, this translates into an increase in the price of domestic goods relative to foreign ones (i.e., a terms of trade effect), which improves the EU trade balance and raises GNE. Please note how results are consistent with [florez2024eu], who report a welfare gain of 0.016%. Despite the increase in GNE, real wages decline by 0.02% and 0.05% under the reduced and full CBAM scenarios, respectively, as a consequence of a higher price index.
Finally, panel (b) registers a moderate effect on GNE in extra-EU countries: $0.008\%$ under the reduced CBAM and $0.02\%$ under the full CBAM scenario. Similar losses in real wages are also recorded in the same table. In this context, concerns about a disproportionate burden on developing or emerging economies seems exaggerated. However, please note that our theoretical setup abstracts from labor market or institutional frictions in the adjustment to the new equilibrium. In this case, it is also more evident that scenario with endogenous prices are not different from the ones with exogenous prices. As already discussed in Section 6.1, our exercises do not consider possible strategic responses by partner countries, which are however an important element of the original CBAM design. The introduction of national carbon markets in extra-EU countries or the investment in less carbon-intensive technologies could help widen the gap between the exogenous and endogenous scenarios we report, and thus help converging to the target of lower emissions.
Variable Reduced CBAM Full CBAM
@@
Table 3: Policy-induced changes in Gross National Expenditure (% of baseline level) as a proxy for welfare. Each value is considered under two scenarios: one in which the CBAM is determined endogenously (Endog.) and one in which it is exogenous (Exog.).
6.4 Separating trade and technology.
@@
Finally, we present two counterfactual exercises to shed light on how technology, on the one hand, and international trade, on the other hand, can influence the effects of CBAM. In the first, we aim to capture the effect of technology; therefore, we vary the output elasticity with respect to emissions. In the second part, we examine the role of the EU’s economic integration within global supply chains. In both exercises, we focus on the key outcome of our analysis — embodied emissions in EU imports.
@@
Let us start with the effect of technology. We construct a counterfactual in which carbon intensities, captured by $\rho_{i}^{j}$, are just half ($50\%$) of those observed at the baseline. This is a circumstance when the world’s technology becomes cleaner. Figure 5 shows the reduction in emissions embodied in EU imports—both direct and total—under the two scenarios. Results are presented as percentage changes relative to the baseline. As expected, the impact of the CBAM on reducing embodied emissions weakens in a world where global production is already cleaner, i.e., less carbon-intensive, since the associated CBAM charges are lower. Interestingly, the reduction is more than proportional, underscoring the disproportionate impact of production technologies on the policy’s effectiveness.
@@
Figure 5: Changes in emissions embodied in EU imports under current and lower carbon intensity are shown with bars. Carbon intensities are lower when they are half of their reference value ($50\%$). Percentages are computed with reference to the value of the corresponding embodied emissions (direct or total) under the scenario with current carbon intensities and no CBAM. Darker colors indicate the amount of emissions in direct imports, while lighter colors indicate the amount of emissions in indirect imports.
@@
Figure 6: Changes in emissions embodied in EU Imports under varying levels of trade integration Percentages are computed with reference to the value of the corresponding embodied emissions (direct or total) under the scenario with the current level of trade integration (0% change) and no CBAM.
Let us now assess the role of international trade, which implies the construction of a counterfactual where we systematically alter the reliance on imported inputs in a range $50\%,+50\%]$. Specifically, for each EU importing sector, we vary the ratio of imported to domestic inputs, where a higher ratio reflects deeper integration of EU production into global supply chains. We then plot the impact of the CBAM on embodied emissions in Figure 6 as a function of the foreign inputs ratio, while holding all else constant. CBAM induces two types of substitution: towards domestic inputs and towards cleaner non-targeted foreign inputs. We find that when the EU economy is less integrated – that is, when sectors rely more on domestic rather than imported inputs – the CBAM leads to a stronger shift toward domestic sourcing. This reallocation results in a larger reduction in emissions embodied in imports, simply because a greater share of inputs originates within the EU and thus lies outside the CBAM scope. As integration increases, the dependence on foreign inputs rises, diminishing the potential for substituting with domestic inputs, and consequently weakening the emissions-reducing effect of the policy. Nevertheless, the CBAM continues to influence sourcing decisions, prompting a shift toward cleaner goods, which partly offsets the reduced scope for domestic substitution. Although emissions embodied in imports still fall relative to the baseline, the overall reduction is smaller, highlighting the dominant role of domestic substitution in driving the policy’s impact.
We conclude that our counterfactuals demonstrate how cleaner global production considerably reduces the scope for CBAM, underscoring the importance of technological progress. On the other hand, a higher level of supply chain integration weakens policy effectiveness, reducing the scope for substitution toward domestic inputs. Eventually, international trade has an impact on CBAM, albeit to a lesser extent than technological progress.
@@ 7. Conclusion
This paper provides a comprehensive assessment of the economic and environmental implications of the EU’s Carbon Border Adjustment Mechanism, utilizing a multi-country, multi-sector general equilibrium model with global input-output linkages. By endogenizing both the CBAM and carbon prices, we capture the feedback mechanisms that arise between national climate policies and international trade patterns. In line with the original policy design, we present a $2\times 2$ alternative scenario to illustrate its implementation. A full and a reduced CBAM scenario are combined with the possibility of endogenizing or not CBAM prices, respectively. Results challenge the notion of a trade-off between eliminating carbon leakage, on the one hand, and gains from welfare and trade, on the other. We find that CBAM crucially induces a reallocation of purchases towards domestic and foreign clean inputs, and significantly reduces emissions embodied in imports. Moreover, a positive albeit small increase is detected for EU Gross National Expenditures (GNE), which is primarily due to a terms-of-trade effect. Finally, counterfactual exercises demonstrate the primary roles of technological progress (i.e., cleaner production) and trade (i.e., increased foreign inputs) in emission reduction, where the first dominates.
Crucially, our findings highlight the importance of supply-chain-wise measures to capture the full carbon footprint in production. Moreover, although it is beyond the scope of this paper to evaluate the feasibility of a CBAM design, our efforts in collecting data have made clear the importance of having an extensive information system capable of collecting and regularly updating data on the emissions generated worldwide by various products. The transitional phase is approaching its end, and it is unclear whether the policy can be in force, as planned, by 2026. More generally, our paper contributes to the study of solutions to international coordination problems when economies are independent and full enforcement cannot be achieved in the face of imminent global challenges, including climate change. We can think of other different international coordination problems (international taxation, trade policy, etc.) that, like environmental regulations, necessitate policies that account for the complex interdependence of global production networks.
@@ A.1 Data
  • Input-Output Tables To calibrate the model on the year 2018, we use the 2021 release of the OECD Inter-Country Input-Output tables [oecd2021], collecting data on expenditures in intermediate inputs $X_{ni}^{kj}$ and final goods $F_{ni}^{j}$, gross output $GO_{i}^{j}$ and value-added $VA_{i}^{j}$. Concerning final consumption, we ignore Gross Fixed Capital Formation and Changes in Inventories and Valuables. Values are reported in millions of U.S. dollars at current prices and the flows of goods and services within and across countries are directly mapped to the cost shares in intermediate goods $\tilde{\omega}_{ni}^{kj}=X_{ni}^{kj}/\sum_{n\in\mathcal{N}}\sum_{j\in\mathcal{J}}X_{ni}^{kj}$ and the shares of final consumption $\alpha_{i}^{j}=\sum_{n\in\mathcal{N}}F_{ni}^{j}/\sum_{n\in\mathcal{N}}\sum_{j\in\mathcal{J}}F_{ni}^{j}$. Finally, the shares of value added are computed as $\beta_{i}^{j}=VA_{i}^{j}/GO_{i}^{j}$. We have a unique country-sector pair (Chile - Motor vehicles, trailers and semi-trailers) with a gross output equal to 0, which can be either due to missing observations or zero production in that sector and country. Since these cases generate computational issues, we input a value of 1 to this observation.
@@
  • Emissions and carbon prices Emissions from production are taken from the OECD Environmental Statistics database. We consider Scope 1 Emissions for each country-sector pair $E_{i}^{j}$, including all direct GHG emissions measured in million tonnes of $\text{CO}_{2}$ equivalent. For the European Union, Iceland and Norway we employ a distinct approach. Emissions data are taken from the European Union Transaction Log (EUTL), recording all transactions carried out under the EU-ETS, with information on each entity covered by the system. The EUTL provides granula, entity-level information on verified emissions and the volume of freely allocated allowances on an annual basis. We aggregate these data at the selected sectoral level and adjust the reported emissions by deducting those covered by freely allocated allowances. To calculate the emissions elasticity, $\rho_{i}^{j}$, these values are multiplied by the corresponding country’s Effective Carbon Rate (ECR), which is then converted into USD using the OECD exchange rate for the year 2018. Effective Carbon Rates represent the total price $t_{i}$ of one ton of $\text{CO}_{2}$ emissions as the sum of carbon taxes, specific taxes on energy use, and the price of tradable emission permits. Emission elasticities are then computed as $\rho_{i}^{j}=t_{i}E_{i}^{j}/GO_{i}^{j}$.